# Vertex Representation of Hyperbolic Tensor Networks

### v1

@@ Abstract
We propose a vertex representation of the tensor network (TN) in the anti-de Sitter space (AdS2+0) that we model on a subset of hyperbolic lattices. The tensors form a network of regular $p$-sided polygons ($p>4$) with the coordination number four. The response to multi-state spin systems on the hyperbolic TN is analyzed for their entire parameter space. We show that entanglement entropy is sensitive to distinguish various hyperbolic geometries whereas other thermodynamic quantities are not. We test the numerical accuracy of vertex TNs in the phase transitions of the first, second, and infinite order at the point of maximal entanglement entropy. The hyperbolic structure of TNs induces non-critical properties in the bulk although boundary conditions significantly affect the total free energy in the thermodynamic limit. Thus developed vertex-type TN can be applied to study low-energy quantum states on AdS.
@@ I Introduction
The classification of numerous phases of matter belongs to one of the fundamental goals in physics. Diverse (numerical) approaches have been developed to identify the phases of matter to sort them into the universality classes. Each numerical approach (method) brings up certain unique features, advancing certain algorithmic properties. However, reaching complex improvements in all aspects cannot be met. Considerable research has been devoted to developing numerical methods to investigate systems in their full intricacy. Alternatively, well-established methods may be additionally modified and further generalized.
Among those methods, the tensor network (TN) techniques have been proven as appropriate candidates for simulating strongly correlated systems. Their popularity still increases and spreads to various fields of physics, such as condensed matter physics TN_top_dual ; Orus2 , high-energy physics HEP1 ; HEP2 , quantum information QI , quantum machine-learning algorithms QML ; QML2 , etc. These techniques describe quantum states by a network of mutually connected tensors. Each tensor represents a vertex where a quantum particle (or a collection of particles) resides. The network of vertices imitates the interaction structure inherently incorporated into the quantum state, typically into the ground state or the lowest-lying ones. The numerical accuracy of tensor connections is controlled by an integer parameter that we call the bond degrees of freedom. The higher the bond degree, the more accurate the quantum state results.
@@
Equivalently to the quantum state, TN can also represent the partition function of a physical system described by a classical Hamiltonian. In this work, we employ this alternative TN approach to evaluate the thermodynamic properties and entanglement entropy focusing on the properties of the underlying lattice geometry. We develop a generalized numerical algorithm based on the density matrix renormalization to analyze the network structure by varying the Gaussian curvature and boundary conditions.
The TN procedure generates large lattices with constant Gaussian negative curvatures on which we study multi-state spin models. We aim to investigate the influence of curved lattices on entanglement entropy (and other thermodynamic quantities). This task relates the famous holographic principle hp1 ; hp2 ; hp3 , the key concept in physics, with the impact on the multidisciplinary research among general relativity, quantum information, and condensed matter physics. The holographic principle states that the volume description of a physical system can be recognized on the system boundary.
The area law alaw of entanglement entropy scales with the surface (boundary) of the volume system in contact with the environment. Since the holographic interpretation has not been fully proved in quantum gravity yet, there has been a persistent need for a deeper investigation of the conjectured equivalency between the anti-de Sitter (AdS) hyperbolic space and the conformal field theories (CFT) on the AdS boundary. Hence, the AdS/CFT correspondence ACC0 ; ACC1 ; ACC2 ; ACC3 ; ACC4 relates the classical description of gravity with the critical system living on the lattice boundary giving rise to the theory of quantum gravity.
As a typical example connecting the holographic principle with the quantum many-body physics in condensed matter (statistical) physics, the renormalization techniques, such as multi-scale entanglement renormalization ansatz MERA , have become essential in numerical studies. This method relates many-body entanglement renormalization using the Ising model with the hyperbolic geometry originating in the entanglement Swingle . This fascinating holographic picture has been fostering intensive research by TN methods Cirac ; TN_top_dual ; Orus ; Banuls . The holographic aspects of the entanglement entropy are actively investigated in the connection of AdS/CFT and TN languages TNAds1 ; TNAds2 ; TNAds3 ; TNAds4 ; TNAds5 .
@@ II Lattice Model
We study a set of lattices classified by a pair of integers ($p,r$) known as the Schläfli symbol Schl . In Fig. 1, we show the Euclidean square lattice and three examples of hyperbolic lattices $(5,4)$, $(6,4)$, and $(15,4)$. We construct the lattices by the regular tessellation of congruent (uniform) polygons. The first integer $p$ describes polygons, such as, e.g., squares ($p=4$), pentagons ($p=5$), or hexagons ($p=6$), etc. The second integer $r$ is the coordination number describing the number of polygons meeting at each vertex (we consider $r=4$ fixed throughout this work; generalization to arbitrary $r$ is straightforward weightpq ).
@@ III Vertex representation
There are two equivalent representations in expressing the partition function. The first representation uses the Boltzmann weights to build up the entire lattice by taking the product of ${\cal W}$ and the sum runs over all $q^{N}$ spin configurations $[\sigma]$, as defined above in Eq. (6). We name it the weight representation. The second representation, namely the vertex representation, expresses the partition function as a product of rank-$r$ tensors ${\cal V}$ where the sum runs over all bond configurations (not the spin ones). Then, the partition function is defined as
@@
$\begin{split}{\cal W}_{\sigma_{i}\sigma_{j}}&=\sum\limits_{a,b=0}^{q-1}U^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}\,D^{\leavevmode\nobreak\ }_{ab}\,U^{\dagger}_{b\,\sigma_{j}}=\sum\limits_{a,b=0}^{q-1}U^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}\,\lambda_{a}^{\leavevmode\nobreak\ }\delta^{\leavevmode\nobreak\ }_{ab}\,U^{\dagger}_{b\,\sigma_{j}}\\ &=\sum\limits_{a=0}^{q-1}\left(U^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}\sqrt{\lambda_{a}^{\leavevmode\nobreak\ }}\right)\left(\sqrt{\lambda_{a}^{\leavevmode\nobreak\ }}\,U^{\dagger}_{a\,\sigma_{j}}\right)=\sum\limits_{a=0}^{q-1}Y^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}Y^{\dagger}_{a\,\sigma_{j}}\end{split}$ (8)
Provided that coupling is ferromagnetic, i.e., $J>0$, the diagonal matrix $D_{ab}=\lambda_{a}\delta_{ab}$ contains only non-negative eigenvalues $\lambda_{a}\geq 0$. Then, the diagonalization can be formulated by multiplying two identical matrices $Y$. Moreover, the unitary matrix $U$ consisting of eigenvectors is also real, yielding $Y^{\dagger}_{\leavevmode\nobreak\ }=Y^{T}_{\leavevmode\nobreak\ }$.
We can now define the tensor
@@
by summing up the $q$-state spin $\sigma=0,\dots,q-1$ (see the graphical visualization in Fig. 3). The rank-$r$ tensor ${\cal V}$ is the desired vertex required in forming the vertex representation of TN. The vertex ${\cal V}$ coincides with the coordination number of the lattice (here, $r=4$). Building up a lattice with an arbitrary constant coordination number $r$ requires preparing a rank-$r$ tensor by summing the product of $r$ matrices $Y$ in Eq. (9). Hence, the weight and vertex representations are equivalent in the sense of ${\cal Z}_{p}^{({\cal W})}={\cal Z}_{p}^{({\cal V})}$ that are interconnected via Eqs. (8) and (9).
The construction of the ($p,4$) lattices follows an algorithm with the iterative scheme. Let the integer $j=1,2,3,\dots$ enumerate the iteration steps. To start, we prepare two initial tensors at the first iteration $j=1$: the rank-3 transfer tensor ${[{\cal P}_{j=1}]}_{abc}\equiv{\cal P}_{abc}$ and the rank-2 corner transfer matrix ${[{\cal C}_{j=1}]}_{ab}\equiv{\cal C}_{ab}$ which are placed on the lattice boundary and define boundary conditions. Figure 4 visualizes the initial construction of ${\cal P}_{abc}$ and ${\cal C}_{ab}$ in analogy to the rank-4 vertex tensor ${\cal V}_{abcd}$. We reserve the integer $k$ to describe the final iteration step in the sequence $j=1,2,3,\dots,k$.
@@ III.1 Recurrent scheme
If we compare the original extension process of the square $(4,4)$ lattice Nis-ver1 ; Nis-ver2 ; Nis-ver3 with the current notation (as visualized in Fig. 5), we can express the recurrence relations in a simplified form $\tilde{\cal C}_{j+1}={\cal V}{\cal P}_{j}^{2}{\cal C}_{j}^{\leavevmode\nobreak\ }$ and $\tilde{\cal P}_{j+1}={\cal V}{\cal P}_{j}^{\leavevmode\nobreak\ }$ on the square lattice. Using this analogy, the expansion of the hyperbolic pentagonal $(5,4)$ lattice obeys the rules $\tilde{\cal C}_{j+1}={\cal V}{\cal P}_{j}^{2}{\cal C}_{j}^{3}$ and $\tilde{\cal P}_{j+1}={\cal V}{\cal P}_{j}^{\leavevmode\nobreak\ }{\cal C}_{j}^{\leavevmode\nobreak\ }$, cf. Fig. 5 and 6.
@@
(There are minor differences between the vertex recurrence relations in Eq. (10) and the weight-type ones weightpq .)
Now we describe how the extended tensors of rank-4 $\tilde{\cal C}_{j+1}$ and rank-5 $\tilde{\cal P}_{j+1}$ are transformed back onto the original rank-2 ${\cal C}_{j+1}$ and rank-3 ${\cal P}_{j+1}$ tensors by renormalization processes. We do this using the isometry $U_{j}$ (originating in the reduced density matrix) which is a unitary projection $U_{j}^{\dagger}U_{j}^{\leavevmode\nobreak\ }=\mathbb{1}$ that reduces the exponentially increasing spin degrees of freedom, i.e.,
@@
This transformation process originates in the density matrix renormalization DMRG1 ; DMRG2 ; DMRG3 . The details are provided in Appendix A.
@@
as the partial trace is taken over the environment degrees of freedom, i.e., all the spins except those on the cut depicted in Fig. 6. The isometry matrices $U_{j}$ are subsequently constructed from the reduced density matrix $\rho^{\prime}_{j}$. The isometries contain selected eigenvectors of $\rho^{\prime}_{j}$ corresponding to the $m$ largest eigenvalues. The integer $m\geq q$ controls the numerical accuracy and is referred to as the bond dimension or the number of states kept DMRG1 ; DMRG2 ; DMRG3 ; DMRGsch .
@@ III.2 Thermodynamic quantities
where $\omega_{i}$ are the $m$ largest eigenvalues. For all hyperbolic lattices ($p>4$), we use the bond number $m$ sufficiently large to assure that the truncation error is as small as $1-\sum_{i=1}^{m}\omega_{i}<10^{-30}$ (within the 128-bit numerical precision).

### v2

@@ Abstract
We propose a vertex representation of the tensor network (TN) for classical spin systems on hyperbolic lattices. The tensors form a network of regular $p$-sided polygons ($p>4$) with the coordination number four. The response to multi-state spin systems on the hyperbolic TN is analyzed for their entire parameter space. We show that entanglement entropy is sensitive to distinguish various hyperbolic geometries whereas other thermodynamic quantities are not. We test the numerical accuracy of vertex TNs in the phase transitions of the first, second, and infinite order at the point of maximal entanglement entropy. The hyperbolic structure of TNs induces non-critical properties in the bulk although boundary conditions significantly affect the total free energy in the thermodynamic limit. Thus developed vertex-type TN can be used for the lowest-energy quantum states on the hyperbolic lattices.
@@ I Introduction
The classification of phases of matter and phase transitions belongs to one of the fundamental goals in physics. Multiple analytic and numerical approaches have been developed to identify the phases of matter and classify them into universality classes. Focusing on numerical methods, we select only those that are suitable for the large number of interacting particles. In this work, we focus on improving the applicability of a numerical method that we have generalized to study models on non-Euclidean lattice geometries. Specifically, we study phase transitions induced by hyperbolic lattice geometry with constant negative Gaussian curvatures.
Among those methods, the tensor network (TN) techniques have been proven as appropriate candidates for simulating strongly correlated systems. Their popularity still increases and spreads to various fields of physics, such as condensed matter physics TN_top_dual ; Orus2 , high-energy physics HEP1 ; HEP2 ; hp1 , quantum information QI , quantum machine-learning algorithms QML ; QML2 , etc. These techniques describe quantum states by a network of mutually connected tensors. Each tensor represents a vertex where a quantum particle (or a collection of particles) resides. The network of vertices imitates the interaction structure inherently incorporated into the quantum state, typically into the ground state or the lowest-lying ones. The numerical accuracy of the tensor connections is controlled by an integer parameter that we call the bond dimension (which denotes the degrees of freedom of the bond). The higher the bond dimension, the more accurately the desired quantum state is calculated.
@@
Equivalently to the quantum state, TN can also represent the partition function of a physical system described by a classical Hamiltonian. In this work, we employ this alternative TN approach to evaluate the thermodynamic properties and entanglement entropy focusing on the properties of the underlying lattice geometry. We develop a generalized numerical algorithm based on density matrix renormalization to analyze the network structure by varying the Gaussian curvature and boundary conditions.
The TN procedure describes a process for creating large hyperbolic lattices with constant negative Gaussian curvatures on which we study multi-state spin models. We aim to investigate the influence of curved lattices on entanglement entropy (and other thermodynamic quantities).
@@ II Lattice Model
We study a set of lattices classified by a pair of integers ($p,r$) known as the Schläfli symbol Schl ; Poincare . In Fig. 1, we show the Euclidean square lattice and three examples of hyperbolic lattices $(5,4)$, $(6,4)$, and $(15,4)$. We construct the lattices by the regular tessellation of congruent (uniform) polygons. The first integer $p$ describes polygons, such as, e.g., squares ($p=4$), pentagons ($p=5$), or hexagons ($p=6$), etc. The second integer $r$ is the coordination number describing the number of polygons meeting at each vertex (we consider $r=4$ fixed throughout this work; generalization to arbitrary $r$ is straightforward weightpq ).
@@ III Vertex representation
There are two equivalent representations in expressing the partition function. The first representation uses the Boltzmann weights to build up the entire lattice by taking the product of ${\cal W}$ and the sum runs over all $q^{N}$ spin configurations $[\sigma]$, as defined above in Eq. (6). We name this the weight representation. The second representation, namely the vertex representation, expresses the partition function as a product of rank-$r$ tensors ${\cal V}$ (defined below) where the sum runs over all bond configurations (not the spin ones). Then, the partition function is defined as
@@
$\begin{split}{\cal W}_{\sigma_{i}\sigma_{j}}&=\sum\limits_{a,b=0}^{q-1}U^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}\,D^{\leavevmode\nobreak\ }_{ab}\,U^{\dagger}_{b\,\sigma_{j}}=\sum\limits_{a,b=0}^{q-1}U^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}\,\lambda_{a}^{\leavevmode\nobreak\ }\delta^{\leavevmode\nobreak\ }_{ab}\,U^{\dagger}_{b\,\sigma_{j}}\\ &=\sum\limits_{a=0}^{q-1}\left(U^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}\sqrt{\lambda_{a}^{\leavevmode\nobreak\ }}\right)\left(U^{\leavevmode\nobreak\ }_{\sigma_{j}\,a}\sqrt{\lambda_{a}^{\leavevmode\nobreak\ }}\right)=\sum\limits_{a=0}^{q-1}Y^{\leavevmode\nobreak\ }_{\sigma_{i}\,a}Y^{\leavevmode\nobreak\ }_{\sigma_{j}\,a}.\end{split}$ (8)
Provided that coupling is ferromagnetic, i.e., $J>0$, the diagonal matrix $D_{ab}=\lambda_{a}\delta_{ab}$ contains only non-negative eigenvalues $\lambda_{a}\geq 0$. Thus, we have decomposed ${\cal W}_{\sigma_{i}\sigma_{j}}$ onto two identical matrices (rank-2 tensors) $Y_{\sigma\,a}=U^{\leavevmode\nobreak\ }_{\sigma\,a}\sqrt{\lambda_{a}^{\leavevmode\nobreak\ }}$ with two distinct variables: the spin variable $\sigma$ and the bond variable $a$ to be used for constructing the entire TN.
Equation (8) describes a symmetric decomposition in the vertex representation sqrt(Q) , where we schematically split the Boltzmann weight ${\cal W}=Y\cdot Y$ into the product of two identical matrices $Y$. The decomposition holds for symmetric Hamiltonians with non-negative eigenvalues of ${\cal W}$. If comparing the decomposition in Eq. (8) with that in Ref. sqrt(Q), where ${\cal W}={\sqrt{\cal W}}\cdot{\sqrt{\cal W}}$, they both become equivalent if $\sqrt{{\cal W}}=Y$ for boundary magnetic field $b=0$.
We intentionally impose $b\neq 0$ on the boundary spins only, see Eq. (3), to induce a faster spontaneous symmetry breaking. This leads to the asymmetric Hamiltonian and there are two ways of the decomposition ${\cal W}=Y\cdot Y^{\prime}$. We use either the eigenvalue decomposition with $Y_{\sigma a}=U_{\sigma a}|d_{a}|$ and $Y^{\prime}_{\sigma a}=\text{sgn}(d_{a})Y_{\sigma a}$ or, equivalently, we can use the singular value decomposition with $Y_{\sigma a}=U_{\sigma a}\sqrt{d_{a}}$ and $Y^{\prime}_{\sigma a}=V_{\sigma a}\sqrt{d_{a}}$ where all singular values are non-negative $d\geq 0$ after decomposing ${\cal W}_{\sigma\sigma^{\prime}}=\sum_{x}U_{\sigma x}d_{x}V_{\sigma^{\prime}x}$.
We can now define the basic vertex tensor
@@
by summing up the $q$-state spin $\sigma=0,\dots,q-1$ (see the graphical visualization in Fig. 3). The rank-$r$ tensor ${\cal V}$ is the desired vertex required in forming the vertex representation of TN. The vertex ${\cal V}$ coincides with the coordination number of the lattice (here, $r=4$). Building up a lattice with an arbitrary constant coordination number $r$ requires preparing a rank-$r$ tensor by summing the product of $r$ matrices $Y$ in Eq. (9). Hence, the weight and vertex representations are equivalent in the sense of ${\cal Z}_{p}^{({\cal W})}={\cal Z}_{p}^{({\cal V})}$ which are interconnected via Eqs. (8) and (9). As specified below, we define two types of tensors: the rank-3 transfer tensor ${\cal P}_{abc}=\sum\limits_{\sigma}Y_{\sigma a}Y_{\sigma b}Y_{\sigma c}$ and the corner transfer matrix ${\cal C}_{ab}=\sum\limits_{\sigma}Y_{\sigma a}Y_{\sigma b}$, as in Fig. 4 in analogy to the rank-4 vertex tensor ${\cal V}_{abcd}$.
The construction of the ($p,4$) lattices follows an algorithm with the iterative scheme. Let the integer $j=1,2,3,\dots$ enumerate the iteration steps. To start, we prepare two initial tensors at the first iteration $j=1$: the rank-3 transfer tensor ${[{\cal P}_{j=1}]}_{abc}\equiv{\cal P}_{abc}$ and the rank-2 corner transfer matrix ${[{\cal C}_{j=1}]}_{ab}\equiv{\cal C}_{ab}$ which are placed on the lattice boundary and define boundary conditions. We reserve the integer $k$ to describe the final iteration step in the sequence $j=1,2,3,\dots,k$.
@@ III.1 Recurrent scheme
If we compare the original extension process of the square $(4,4)$ lattice Nis-ver1 ; Nis-ver2 ; Nis-ver3 with the current notation (as visualized in Fig. 5), we can express the recurrence relations in a simplified form $\tilde{\cal C}_{j+1}={\cal V}{\cal P}_{j}^{2}{\cal C}_{j}^{\leavevmode\nobreak\ }$ and $\tilde{\cal P}_{j+1}={\cal V}{\cal P}_{j}^{\leavevmode\nobreak\ }$ on the square lattice. Using this analogy, the expansion of the hyperbolic pentagonal $(5,4)$ lattice obeys the rules $\tilde{\cal C}_{j+1}={\cal V}{\cal P}_{j}^{2}{\cal C}_{j}^{3}$ and $\tilde{\cal P}_{j+1}={\cal V}{\cal P}_{j}^{\leavevmode\nobreak\ }{\cal C}_{j}^{\leavevmode\nobreak\ }$, cf. Fig. 5.
@@
These recurrence relations of TN are expressed concisely in Eq. (10). They represent tensor contractions, as specified in Appendix A. (There are minor differences between the vertex recurrence relations in Eq. (10) and the weight-type ones weightpq .)
Now we describe how the extended tensors of rank-4 $\tilde{\cal C}_{j+1}$ and rank-5 $\tilde{\cal P}_{j+1}$ are transformed back onto the original rank-2 ${\cal C}_{j+1}$ and rank-3 ${\cal P}_{j+1}$ tensors by renormalization processes. We do this using the isometry $U_{j+1}$ (originating in the reduced density matrix) which is a unitary projection $U_{j+1}^{\dagger}U_{j+1}^{\leavevmode\nobreak\ }=\mathbb{1}$ that reduces the exponentially increasing spin degrees of freedom, i. e.,
@@
This transformation process originates in the density matrix renormalization DMRG1 ; DMRG2 ; DMRG3 . The computational cost of the optimized CTMRG code is $O(pq^{3}m^{3})$ only. The details are provided in Appendix A.
@@
as the partial trace is taken over the environment degrees of freedom, i.e., all the spins except those on the cut depicted in Fig. 6. The isometry matrices $U_{j+1}$ are subsequently constructed from the reduced density matrix $\rho^{\prime}_{j+1}$. The isometries contain selected eigenvectors of $\rho^{\prime}_{j+1}$ corresponding to the $m$ largest eigenvalues. The integer $m\geq q$ controls the numerical accuracy and is referred to as the bond dimension or the number of states kept DMRG1 ; DMRG2 ; DMRG3 ; DMRGsch .
A remark on the Hermiticity of the reduced density matrix $\rho^{\prime}$ is needed at this stage. For hyperbolic lattices $(p,4)$, $\rho^{\prime}$ is Hermitian for even integers $p=4,6,8,\dots$ From Fig. 6, we can see that the lattice can be divided into two identical halves only if $p$ is even, and each half of the lattice consists of $p/2$ corner transfer matrices. Their partial sum approximates the state $|\Psi\rangle=\sum^{\prime}\tilde{\cal C}_{\leavevmode\nobreak\ }^{p/2}$ so that
$\rho^{\prime}={\rm Tr}^{\prime}(\tilde{\cal C}_{\leavevmode\nobreak\ }^{p})={\rm Tr}^{\prime}(\tilde{\cal C}_{\leavevmode\nobreak\ }^{p/2}\tilde{\cal C}_{\leavevmode\nobreak\ }^{p/2})={\rm Tr}^{\prime}|\Psi\rangle\langle\Psi|.$ (13)
We keep the state normalized, i.e., $\langle\Psi|\Psi\rangle=1$, for evaluating entanglement entropy correctly.
The odd values of $p$, lead to $\rho^{\prime}$ that is non-Hermitian, and we use such construction that involves the least asymmetry (and demand normalization $\langle\Phi|\Psi\rangle=1$), i.e.,
$\rho^{\prime}={\rm Tr}^{\prime}(\tilde{\cal C}_{\leavevmode\nobreak\ }^{p})={\rm Tr}^{\prime}(\tilde{\cal C}_{\leavevmode\nobreak\ }^{(p-1)/2}\tilde{\cal C}_{\leavevmode\nobreak\ }^{(p+1)/2})={\rm Tr}^{\prime}|\Psi\rangle\langle\Phi|.$ (14)
In Ref. triangular, , we studied hyperbolic lattices with non-Hermitian $\rho^{\prime}$ implementing three distinct approaches: (1) diagonalization of non-symmetric matrices, (2) singular value decomposition, and (3) diagonalization of symmetrized $\rho^{\prime}_{\rm sym}$. We encountered numerical instabilities in the cases (1) and (2), which prevent the algorithm to converge. In case (3), the diagonalization of the symmetrized reduced density matrix $\rho^{\prime}_{\rm sym}=\tfrac{1}{2}(|\Psi\rangle\langle\Phi|+|\Phi\rangle\langle\Psi)|)$ always resulted in numerically stable solutions. This transform yields the optimal isometries $U$ with systematic and stable behavior of the algorithm.
All numerical quantities, including the correct phase transition temperatures of the current vertex TN completely coincide with the weight representation of TN weightp4 ; weightpq where the hyperbolic lattices $(p,4)$ always lead to the Hermitian (symmetric) construction of $\rho^{\prime}$. Alternatively, discovering a dual set of bi-orthonormal bases (corresponding to the left and right eigenvectors) can be a possible choice to reproduce the current results BTMRG .
@@ III.2 Thermodynamic quantities
where we consider the $m$ largest eigenvalues $\omega_{i}$. For all hyperbolic lattices ($p>4$), we use a sufficiently large bond number $m$ to ensure that the truncation error is as small as $1-\sum_{i=1}^{m}\omega_{i}<10^{-30}$ (within the 128-bit numerical precision).
An additional remark on quantum entanglement entropy for classical spin systems is in place. We can ascribe classical spin models to quantum spin models using a quantum-classical correspondence QCC ; Chatelain . The correspondence maps a $D$-dimensional quantum system onto a $(D+1)$-dimensional classical one due to the equivalence of their partition functions. The extra dimension in classical systems originates in the imaginary-time evolution via the Suzuki-Trotter expansion Trotter ; Suzuki1 ; Suzuki2 .
In summary, the classical spin models in this work (studied on various $2D$ lattice surfaces) could be related to $1D$ quantum spin systems. For instance, the von Neumann entanglement entropy computed for the classical Ising model is identical to the associated quantum Ising model (after rescaling the transverse magnetic field $h\to h/h_{\rm c}$ and temperature $T\to T/T_{\rm c}$ in the corresponding classical model).